OpenStudy (vera_ewing):

If a hospital patient is given 50 milligrams of medicine which leaves the bloodstream at 10% per hour, how many milligrams of medicine will remain in the system after 6 hours? Use the function A(t) = Iert. (1 point) 12.39 mg 27.44 mg 45.24 mg 91.11 mg

zepdrix (zepdrix):

\[\Large\rm A(\color{orangered}{t})=\mathcal I e^{r\color{orangered}{t}}\]Ok, so I represents the `initial amount`. Looks like the patient was given 50 to start, yes? :)\[\Large\rm A(\color{orangered}{t})=50 e^{r\color{orangered}{t}}\]

OpenStudy (vera_ewing):

Ok, now what?

zepdrix (zepdrix):

Mmm let's make sure we're understanding what the rate, \(\Large\rm r\), should be. Hmm

zepdrix (zepdrix):

If 10% is leaving the bloodstream each hour, then 90% will remain, yah?

OpenStudy (vera_ewing):

Ok, yes.

zepdrix (zepdrix):

Remember how to write 90% as a decimal? We want it as a decimal when we toss it into the equation.

OpenStudy (vera_ewing):

0.9

zepdrix (zepdrix):

\[\Large\rm A(\color{orangered}{t})=50 e^{0.9\cdot\color{orangered}{t}}\]Ok good.

zepdrix (zepdrix):

So t represents time, more specifically, hours. So they want us to calculate the amount of medicine in the patients system at time t=6 hours.

zepdrix (zepdrix):

\[\Large\rm A(\color{orangered}{6})=50 e^{0.9\cdot\color{orangered}{6}}\]

zepdrix (zepdrix):

And then we let the calculator do the work from there :)

zepdrix (zepdrix):

Ooo that did not give a nice calculator answer :3 sec lemme check that again..

zepdrix (zepdrix):

Mmm I guess I screwed up the rate :( Or the question was worded poorly, one of those two things lol. I'm not sure >.< So we were supposed to use the 10%, not the 90%.\[\Large\rm A(\color{orangered}{6})=50 e^{0.1\cdot\color{orangered}{6}}\]

zepdrix (zepdrix):

Do you have a calculator?

OpenStudy (vera_ewing):

Yeah, I can use mathway.com , but I don't know how to input it. Can you do it?

zepdrix (zepdrix):

Put it into the calculator something like this,\[\Large 50\times e\text{^}(0.1\times6)\]

OpenStudy (vera_ewing):

I can't input that for some reason :(

zepdrix (zepdrix):

In mathway it should look like this:\[\Large\rm 50\cdot e^{(0.1\cdot6)}\]

zepdrix (zepdrix):

No why not? D:

OpenStudy (vera_ewing):

ok so I got 91.10

zepdrix (zepdrix):

Ok good! :) Make sure you round to the hundredth place correctly though. That's a 5 as the next number ya?

zepdrix (zepdrix):

91.105

OpenStudy (vera_ewing):

right, so 91.11 :)

OpenStudy (vera_ewing):

I got it wrong!

zepdrix (zepdrix):

Was not D? D:

OpenStudy (vera_ewing):

No it wasn't D...that's so weird!

zepdrix (zepdrix):

Wait wait why would it be D, if he started with 50 milligrams and its decreasing, how does he end up with 90 later on. omggggg im stupid today >.<

zepdrix (zepdrix):

*cri*

OpenStudy (vera_ewing):

Omg so what is it?

zepdrix (zepdrix):

\[\Large\rm A(\color{orangered}{6})=50 e^{(-0.1\cdot\color{orangered}{6})}\]Negative 10%, 10 percent leaving the bloodstream. Which works out to 27.440581 mg. I think that's the right answer. Ahhh I'll have to brush up on my exponential decay problems :( ugh

OpenStudy (vera_ewing):

oh ok. Thank you